### v1

# Understanding via Gaze: Gaze-based Task Decomposition for Imitation Learning of Robot Manipulation

Abstract
In imitation learning for robotic manipulation, decomposing object manipulation tasks into multiple semantic actions is essential. This decomposition enables the reuse of learned skills in varying contexts and the combination of acquired skills to perform novel tasks, rather than merely replicating demonstrated motions. Gaze, an evolutionary tool for understanding ongoing events, plays a critical role in human object manipulation, where it strongly correlates with motion planning. In this study, we propose a simple yet robust task decomposition method based on gaze transitions. We hypothesize that an imitation agent’s gaze control—fixating on specific landmarks and transitioning between them—naturally segments demonstrated manipulations into sub-tasks. Notably, our method achieves consistent task decomposition across all demonstrations, which is desirable in contexts such as machine learning. Using teleoperation—a common modality in imitation learning for robotic manipulation—we collected demonstration data for various tasks, applied our segmentation method, and evaluated the characteristics and consistency of the resulting sub-tasks. Furthermore, through extensive testing across a wide range of hyperparameter variations, we demonstrated that the proposed method possesses the robustness necessary for application to different robotic systems.
I INTRODUCTION
Imitation stands as a uniquely promising approach for enabling robots to function as autonomous agents capable of acquiring object manipulation skills that align with human intentions. In fact, not only robots but also humans learn how they can and should handle various objects by observing expert demonstrations—such as young children learning from adults, or novices learning from skilled practitioners [1]. However, simply replicating demonstrated motions does not allow an agent to reuse learned skills in slightly altered contexts, nor to combine previously acquired motions into novel object manipulations.
To move beyond such a rote imitation, it is crucial that imitative agents gain a deeper understanding of demonstrated object manipulations, which requires autonomously segmenting a sequence of demonstrated motions into semantically meaningful units. The importance of such uniform task decomposition as the basis of action understanding is suggested by human studies—such as those by Newtson et al.—showing remarkable consistency among observers when segmenting video-recorded human actions [2].
Figure 1: Gaze itself possesses an inherent structure in object manipulation, which enables task decomposition by detecting transitions between gaze landmarks during imitation. (a) Gaze transition of a human teleoperating a robot. (b) Gaze transition of a human making a sandwich [3]. (c) Overview of our proposed method.
How, then, can we achieve autonomous and practical task decomposition of object manipulations? As Yarbus famously stated, “Eye movement reflects the human thought processes; so the observer’s thought may be followed to some extent from records of eye movement,” suggesting that eye gaze serves as a naturally evolved tool for understanding ongoing events [4]. In particular, it is well-established that during human object manipulations, individuals repeatedly fixate their gaze on task-relevant landmarks for certain periods and then move on to the next (Fig. 1), and that these gaze transition behaviors are tightly coupled with motor planning processes [5, 6, 7]. Furthermore, the work of Fathi et al. demonstrates that visual features located in the vicinity of a subject’s gaze during daily activities correlate with the verbal descriptions of these actions [3].
Taken together, these insights suggest that focusing on gaze position and the associated visual features offers a pathway to understanding—and thus segmenting—object manipulations. In this paper, we show that a simple approach of detecting transitions between gaze landmarks during imitation enables autonomous, general, and robust task decomposition of object manipulation sequences (Fig. 1). Here, although it would be preferable for the robot to autonomously control its own gaze during imitation, acquiring a general-purpose gaze-control policy remains challenging. Therefore, following the method proposed in [8], we measure the expert’s gaze during imitation (Fig. 1) and use it as a stand-in for the robot’s gaze. In other words, we assume an ideal scenario of joint attention, where the imitating agent and the human demonstrator share identical gaze transitions throughout the imitation process. This substitution is made feasible by the fact that the expert demonstrations are generated through robot teleoperation by a human. Under teleoperation, object manipulation typically relies solely on visual feedback, placing the human teleoperator under the same conditions as the imitating agent, which must interpret visual information to understand the demonstrated object manipulations.
In summary, the main contributions of this paper are follows:
  • We propose a practical task decomposition method that decomposes object manipulations in demonstrations into sub-tasks in a simple yet robust manner, by detecting the teleoperator’s gaze transitions during demonstrations.
  • We show that an imitation agent’s gaze control can simultaneously achieves task decomposition, underscoring the importance of developing autonomous gaze control capabilities in robotic object manipulation.
II RELATED WORK
II-A Robot Manipulation Imitation Using Task Decomposition
Many approaches have been proposed that utilize temporal task decomposition for imitation learning in object manipulation tasks. A common strategy for enabling robustness to long-horizon tasks is to employ goal-conditioned policy models [9, 10]. Furthermore, recent years have seen an increase in language-conditioned models aimed at multi-task capabilities [11, 12, 13, 14]. These goal- or language-conditioned methods implicitly assume that object manipulations in the given demonstrations are segmented into individual sub-tasks, and it has also been reported that providing language instructions per sub-task during inference improves task success rates [15].
However, in most existing work, the segmentation of these demonstrations is performed manually. Automating this process, as proposed in our research, can enhance robot autonomy and reduce annotation costs.
II-B Conventional Task Decomposition Methods
A variety of methods have been proposed for temporally segmenting actions—including object manipulation—over time. Nonetheless, few are suited specifically for segmenting demonstrations of robot object manipulation.
Among the existing methods are those that segment actions based on hand or body trajectories and velocities [16, 17, 18]. Such methods require body tracking and are not robust to trajectory variations. In contrast, our proposed approach does not rely on directly measuring or leveraging the motion itself. Other methods detect segmentation points through changes in the relationship between objects and the hand [19, 20], but they demand advanced image processing techniques, such as object or hand segmentation, detection, and tracking. Although further investigation is required to determine which type of segmentation is more beneficial for imitation learning, there is a qualitative difference between the segmentation results of these methods and those produced by our approach. Notably, our method does not necessarily cut actions at the moment of contact.
With the advent of deep learning, there has been a rise in end-to-end approaches that take images as input and directly output action labels [21, 22, 23, 24]. However, applying these methods to previously unseen tasks requires additional data and annotations, and they do not guarantee uniform segmentation. By contrast, our method can perform task decomposition without any supervision, as long as gaze data is available. Although there have been a few studies focusing on gaze—particularly on image features around the gaze position [25, 3]—our approach explicitly ties the timing of gaze transitions to the timing of sub-task changes, achieving a simple and uniform decomposition.
III METHOD
In this section, we first formulate task decomposition of object manipulation based on gaze transitions. We then detail our implementation procedure. Specifically, we begin by explaining how gaze transitions are detected from time-series gaze data, and subsequently describe the refinement process used to achieve consistent segmentation across the demonstration dataset.
III-A Gaze-based Task Decomposition
We consider a scenario in which an agent that has acquired gaze control imitates object manipulation using only visual information. Let $\{g_{t}\}_{t=0}^{T}$ denote the time-series of the agent’s gaze positions, and $\{f_{t}\}_{t=0}^{T}$ the time-series of the visual features observed around those gaze positions.
When the object manipulation is decomposed into a set of sub-tasks $\{\tau_{k}\}_{k=0}^{K}$, the objective in this section is to find the $K-1$ time steps at which the sub-tasks switch (i.e. $\{t\mid k(t+1)=k(t)+1\}$).
In human object manipulation, it is well-established that each task has a specific set of gaze landmarks $\{c_{i}\}_{i=0}^{N}$, and that humans fix their gaze on each landmark for a certain duration, then sequentially transition to the next landmark [6, 7, 5]. Here, the landmark index $i$ is arranged according to the order in which landmarks are fixated from the start of the task (duplicates may occur).
In this study, we hypothesize that the time steps at which the index of the fixated gaze landmark changes, $\{t\mid i(t+1)=i(t)+1\}$, corresponds to the time steps at which the sub-tasks switch, $\{t\mid k(t+1)=k(t)+1\}$. In other words, we realize task decomposition by detecting transitions of gaze landmarks.
Let $p^{i(t)}_{t}$ be the location of the landmark $c_{i(t)}$ in the agent’s field of view at time step $t$. Then, the gaze position $g_{t}$ and the visual feature $f_{t}$ at that gaze position can be written as follows:
$g_{t}=p\bigl{(}c_{i=i(t)},t\bigr{)}+\epsilon_{t},$ (1)
$f_{t}=f\bigl{(}g_{t}\bigr{)}+\zeta_{t}=f\bigl{(}p(c_{i=i(t)},t)+\epsilon_{t}\bigr{)}+\zeta_{t}.$ (2)
where $\epsilon_{t}$ and $\zeta_{t}$ represent noise in the gaze position and perturbations in the visual features caused, for example, by hand movements, respectively, and $f(\cdot)$ is the visual feature extractor at the given gaze position.
In other words, by leveraging the relationships in Eqs. (1) and (1), we can detect the transitions of gaze landmarks from the gaze information $\{g_{t}\}$ and $\{f_{t}\}$, and thereby achieve task decomposition of the demonstrated object manipulation.
III-B Gaze Transition Detection
We now describe the implementation details for segmenting demonstration tasks based on gaze transitions. As discussed in the previous section, an algorithm for detecting the time steps at which gaze landmarks switch, using time-series gaze data, is critical for task decomposition.
III-B1 Median Filtering
Gaze data typically contain noise. Even when the agent is fixating on a stationary gaze landmark, the recorded gaze positions often exhibit fluctuations. Moreover, the agent may occasionally shift its gaze for a short period to a different landmark during object manipulation. To handle such noise in a simple yet effective manner, we adopted a median-filtering approach. Specifically, for the agent’s gaze position $g_{t}$ we applied the process shown in Eq. (3) with a window size $w$ as a hyperparameter, and used the filtered gaze data $\hat{g}_{t}$ in all subsequent processing:
$\hat{g}_{t}=\text{median}\bigl{(}g_{t-\tfrac{w}{2}:t+\tfrac{w}{2}}\bigr{)}.$ (3)
III-B2 Visual Features Around the Gaze
Gaze transitions cause not only positional changes in the gaze but also variations in the visual features around the gaze. To detect these transitions from visual features, it is necessary to employ a feature extractor based on high-quality representation learning. For each time step, we cropped an $b\times b$ region centered at the filtered gaze position $\hat{g}_{t}$ from both the left and right images and fed these cropped images into the pre-trained CLIP model [26] to extract their feature vectors $f_{t}\in\mathbb{R}^{512\times 2}$. This time-series of feature vectors, along with the gaze position, was used for gaze transition detection.
III-B3 Change Detection
For the obtained gaze position $\hat{g}_{t}$ and visual feature $f_{t}$, we computed the difference from the previous time step to detect abrupt changes in the time-series data caused by gaze transitions:
$s^{pos}_{t}=\bigl{\lVert}\hat{g}_{t}-\hat{g}_{t-1}\bigr{\rVert},$ (4)
$s^{feat}_{t}=-\frac{\log\bigl{(}{f^{left}_{t-1}}^{T}f^{left}_{t}+1\bigr{)}+\log\bigl{(}{f^{right}_{t-1}}^{T}f^{right}_{t}+1\bigr{)}}{2}$
$+\log 2.$ (5)
Here, $s^{pos}_{t}$ and $s^{feat}_{t}$ represent the change scores for the gaze position and the visual features, respectively. Specifically, $s^{pos}_{t}$ is the Euclidean distance between the two 4-dimensional vectors representing the left and right gaze positions. Meanwhile, $s^{feat}_{t}$ is derived from the inner products between the corresponding gaze feature vectors in consecutive time steps, normalized to ensure non-negativity, and then averaged for the left and right features.
We introduced the thresholds $\theta_{pos}$ and $\theta_{feat}$ as hyperparameters for these two change scores, and identified any time step at which both change scores exceeded their respective thresholds as a change point $d_{i}\in D$.
III-C Refinement Based on Segmentation Uniformity
From a machine learning perspective—particularly one that relies on multiple demonstrations—it is desirable for all demonstrations of the same task to undergo a uniform task decomposition. For example, in a pick-and-place task, mixing one demonstration segmented into “pick” and “place” with another in which these two motions are grouped into a single segment is not ideal for building a consistent training dataset. Therefore, in this study, we introduce a refinement procedure, $D\rightarrow D^{*}$, for each demonstration to ensure that demonstrations of the same task share as consistent a number of sub-tasks as possible.
First, we enumerate the number of change points $\lvert D\rvert$ obtained via the gaze transition detection (Section III-B) for all demonstrations and denote the most frequent value by $s$. We assume $s$ to be the number of sub-tasks for the task and adjust the thresholds $\theta_{pos}$ and $\theta_{feat}$ so that the number of segments in each demonstration matches $s$. Specifically, for any demonstration whose segmented count is smaller than $s$, we reduce both thresholds by 1% at each iteration; conversely, for a count larger than $s$, we increase them by 1%. If no threshold values are found that yield $s$ segments within a maximum number of iterations, that demonstration is treated as a detection failure and excluded from the training data.
Finally, the refined set of change points $D^{*}$ obtained by this process is taken as the sub-task boundary for each demonstration, enabling us to divide as many demonstrations as possible into a consistent number of sub-tasks.
IV EXPERIMENTS
IV-A Robot System
Our group previously developed a dual-arm robotic system designed for imitation learning through teleoperation [8]. A human teleoperator controls the robot remotely by viewing its environment through a head-mounted display (HMD), which simultaneously records the operator’s eye gaze. The system can be configured to use either a physical dual-arm robot (a pair of UR5 arms) or a simulated counterpart, enabling the collection of expert demonstrations in both real and virtual settings [27]. The recorded gaze data are provided as pixel coordinates corresponding to the images displayed on the HMD.
In this study, we evaluated our task decomposition method using an expert demonstration dataset collected by the authors. Each demonstration includes synchronized left and right image streams, along with corresponding gaze data, and all of these were employed in our experiments.
IV-B Task Setup
We evaluated the proposed methods on three object manipulation tasks, two conducted using the physical robot system and one within a simulated environment.
We first collected approximately 100 demonstrations for each task as follows:
  • WrapCandy (real robot): In this task, a green candy and a brown handkerchief are arranged on a table with a green tablecloth. The position and orientation of the two objects are randomized on the table. The robot grasps the candy and place it in the middle of the handkerchief. The robot then picks up the (top) left corner of the handkerchief and folds it to wrap the candy.
  • PenInCup (real robot): In this task, a blue marker pen and a red cup are placed on a table; the position and orientation of the two objects is randomized on the table. The robot picks up the pen and puts it in the cup.
  • PileBox (simulator robot): In this task, red and blue boxes are placed on a wooden table; the position and orientation of the two boxes is randomized. The robot grabs the red box and places it on top of the green box.
IV-C Task Decomposition Results
For each task, we performed task decomposition on all demonstrations in the dataset. In doing so, we used the same hyperparameters across all demonstrations and tasks: $w=20,b=256\;(\text{overall image size: }1280\times 720),\theta_{pos}=50,\theta_{feat}=0.05$. Fig. 2 presents an representative example of segmentation results for each task.
Figure 2: Typical task decomposition results using the proposed method for the tasks of WrapCandy (a), PenInCup (b), and PileBox (c). The unshaded square areas represent the images cropped around the gaze positions.
In the WrapCandy task, the series of object manipulations was segmented into four actions: (1) picking up the candy, (2) placing the candy at the center of the towel, (3) grasping the upper-left corner of the towel, and (4) folding the towel. In the PenInCup task, the demonstration was segmented into two actions: (1) picking up the pen and (2) inserting the pen vertically into the cup. In the PileBox task, the demonstration was similarly segmented into two actions: (1) picking up the red box and (2) placing it on top of the green box. We can see that none of these decompositions deviate from human intuition.
A closer examination of the states before and after each segmentation revealed two common features across the tasks:
  • Each sub-task began when the agent (released the previous object and) started reaching for the next relevant object.
  • Each sub-task ended once the action’s completion was visually confirmed. (In a grasping action, for example, the sub-task did not end when the hand merely contacted the object, but rather when the object was lifted slightly off the table, confirming a successful grasp.)
Crucial aspects of object manipulation exist even before actual contact with the object. For example, how the agent reaches for the object significantly influences the success of the manipulation. For this reason, non-contact actions such as reaching should also be included in the segment. However, methods that rely solely on outcome-based cues, such as contact events, cannot detect these reaching actions.
IV-D Uniformity of Task Decomposition
From a practical perspective, it is desirable that all demonstrations within each task exhibit a uniform pattern of task decomposition. Table I shows the number of demonstrations classified as “minority data,” in which the segmentation timings or the number of sub-tasks differed from the most common (majority) decomposition for each task.
In addition, we also evaluated decomposition uniformity under the following ablation settings: (1) without the refinement process (w/o refine), (2) detecting changes based solely on the gaze position $\hat{g}_{t}$ (w/o feat), (3) detecting changes based solely on visual features around the gaze $f_{t}$ (w/o pos), and (4) omitting the refinement process from these two settings (w/o feat, refine and w/o pos, refine).
TABLE I: Number of minority demonstrations exhibiting task decomposition with different timing or segment counts from the majority in each task.
WrapCandy PenInCup PileBox Total
Ours 0 / 64 0 / 125 0 / 147 0 / 336
- w/o refine 1 1 3 5
- w/o feat 0 0 0 0
- w/o feat, refine 1 1 3 5
- w/o pos 0 0 0 0
- w/o pos, refine 4 4 45 53
When the refinement process was applied, we found that all demonstrations across all tasks were segmented in a consistent manner. In contrast, removing the refinement process led to some demonstrations whose decomposition differed from the majority. The number of such minority cases increased noticeably in settings where the gaze position was not used (i.e., w/o pos, refine) compared to those where it was used (i.e., w/o refine and w/o feat, refine). In each scenario, we observed instances of over-segmentation when the gaze shifted slightly during an action, or missed segmentation when the gaze displacement was too small (e.g., when objects were close together). Furthermore, in the condition where only gaze features were employed, change detection was triggered whenever the hand or an object grasped by the hand entered the cropped image region, resulting in over-segmentation.
Figure 3: Number of successfully decomposed demonstrations when varying the hyperparameters. (a) Without refinement process. (b) With refinement process.
IV-E Hyperparameter Robustness
Ensuring that the task decomposition results remain consistent despite changes in hyperparameters is critical for the generality and robustness of the proposed method. In this section, we investigate how variations in the hyperparameters $w,b,\theta_{pos},\theta_{feat}$ used by our method affect the number of correctly segmented demonstrations.
In the previous experiment (Table I), we confirmed that, when gaze position was employed for change detection, the presence or absence of gaze feature data had no significant impact on the segmentation results. Therefore, in this experiment, we performed task decomposition using only the gaze position. Under this condition, the relevant hyperparameters are $w$ and $\theta_{pos}$. Specifically, $w$ determines the filtering strength of the median filter applied to the gaze position, while $\theta_{pos}$ determines the sensitivity for detecting gaze transitions.
Figure 3 shows the number of demonstrations that were correctly segmented when $w$ was set to $\{5,10,15,20,25,30\}$ and $\theta_{pos}$ was set to $\{25,50,75,100,125,150\}$.
When hyperparameters varied extensively, the number of demonstrations that resulted in incorrect decomposition exhibited gradual fluctuations in the absence of refinement processing. In contrast, with refinement processing applied, decomposition was successfully achieved for all demonstrations across the broad parameter range. These findings allow us to conclude that the proposed method possesses the robustness necessary for application to various robotic systems.
In most cases, incorrect decomposition persisted after the refinement process (including instances where the number of successfully segmented demonstrations decreased post-refinement) because decomposition errors resulting from an incorrect number of segments had already become predominant beforehand. Consequently, instead of achieving its intended purpose, the refinement process altered demonstrations that had been correctly segmented initially, leading to an incorrect number of sub-tasks.
Furthermore, setting the window width $w$ to 25 or higher caused the influence of median filtering to become excessively strong. As a result, particularly in the WrapCandy task, it became impossible to detect the required number of segmentation points during the refinement process, regardless of how much the threshold $\theta_{pos}$ was lowered.
V CONCLUSIONS
In this study, we proposed a robust and generalizable method for the autonomous task decomposition of object manipulation behaviors, which serves as a foundation for enabling robots, as autonomous agents, to move beyond rote imitation and to understand appropriately demonstrated object manipulations. This method achieves segmentation solely through detecting the agent’s gaze transitions. Notably, the incorporation of a refinement process allows for consistent task decomposition, which is particularly desirable in applications such as machine learning. Our findings suggest that gaze plays a crucial role in cognition and learning during object manipulation for autonomous agents, including humans. Once a robot acquires gaze control, it can autonomously perform human-like segmentation, which serves as the basis for diverse learning processes, relying solely on the observation of its own gaze transitions.
To achieve greater autonomy in the future, robots must be capable of directly observing and imitating human behaviors rather than relying on teleoperation. In such situations, even if human gaze data were available, using it as a substitute for the robot’s gaze would be inappropriate because humans utilize other sensory modalities, such as tactile information, in addition to visual input for object manipulation. Therefore, breakthroughs in autonomous gaze control are urgently needed today, either for such cases or for cases where demonstration datasets simply lack gaze data from teleoperators, in order to make robots more intelligent.

### v2

@@
-Understanding via Gaze: Gaze-based Task Decomposition for Imitation Learning of Robot Manipulation
+Gaze-Based Task Decomposition for Robot Manipulation in Imitation Learning
 Abstract
-In imitation learning for robotic manipulation, decomposing object manipulation tasks into multiple semantic actions is essential.
+In imitation learning for robotic manipulation, decomposing object manipulation tasks into multiple sub-tasks is essential.
 This decomposition enables the reuse of learned skills in varying contexts and the combination of acquired skills to perform novel tasks, rather than merely replicating demonstrated motions.
-Gaze, an evolutionary tool for understanding ongoing events, plays a critical role in human object manipulation, where it strongly correlates with motion planning.
+Gaze plays a critical role in human object manipulation, where it is strongly correlated with hand movements.
+We hypothesize that an imitating agent’s gaze control—fixating on specific landmarks and transitioning between them—simultaneously segments demonstrated manipulations into sub-tasks.
 In this study, we propose a simple yet robust task decomposition method based on gaze transitions.
-We hypothesize that an imitation agent’s gaze control—fixating on specific landmarks and transitioning between them—naturally segments demonstrated manipulations into sub-tasks.
-Notably, our method achieves consistent task decomposition across all demonstrations, which is desirable in contexts such as machine learning.
-Using teleoperation—a common modality in imitation learning for robotic manipulation—we collected demonstration data for various tasks, applied our segmentation method, and evaluated the characteristics and consistency of the resulting sub-tasks.
+The method leverages teleoperation, a common modality in robotic manipulation for collecting demonstrations, in which a human operator’s gaze is measured and used for task decomposition as a substitute for an imitating agent’s gaze.
+Notably, our method achieves consistent task decomposition across all demonstrations for each task, which is desirable in contexts such as machine learning.
+We applied this method to demonstrations of various tasks and evaluated the characteristics and consistency of the resulting sub-tasks.
 Furthermore, through extensive testing across a wide range of hyperparameter variations, we demonstrated that the proposed method possesses the robustness necessary for application to different robotic systems.
 I INTRODUCTION
+Figure 1: Gaze itself possesses an inherent structure in object manipulation, which enables task decomposition by detecting transitions between gaze landmarks during demonstration.
+(a) Gaze transition of a human teleoperating a robot.
+(b) Overview of our proposed method.
 Imitation stands as a uniquely promising approach for enabling robots to function as autonomous agents capable of acquiring object manipulation skills that align with human intentions.
-In fact, not only robots but also humans learn how they can and should handle various objects by observing expert demonstrations—such as young children learning from adults, or novices learning from skilled practitioners [1].
-However, simply replicating demonstrated motions does not allow an agent to reuse learned skills in slightly altered contexts, nor to combine previously acquired motions into novel object manipulations.
-To move beyond such a rote imitation, it is crucial that imitative agents gain a deeper understanding of demonstrated object manipulations, which requires autonomously segmenting a sequence of demonstrated motions into semantically meaningful units.
-The importance of such uniform task decomposition as the basis of action understanding is suggested by human studies—such as those by Newtson et al.—showing remarkable consistency among observers when segmenting video-recorded human actions [2].
-Figure 1: Gaze itself possesses an inherent structure in object manipulation, which enables task decomposition by detecting transitions between gaze landmarks during imitation.
-(a) Gaze transition of a human teleoperating a robot.
-(b) Gaze transition of a human making a sandwich [3].
-(c) Overview of our proposed method.
-How, then, can we achieve autonomous and practical task decomposition of object manipulations?
-As Yarbus famously stated, “Eye movement reflects the human thought processes; so the observer’s thought may be followed to some extent from records of eye movement,” suggesting that eye gaze serves as a naturally evolved tool for understanding ongoing events [4].
-In particular, it is well-established that during human object manipulations, individuals repeatedly fixate their gaze on task-relevant landmarks for certain periods and then move on to the next (Fig.
-1), and that these gaze transition behaviors are tightly coupled with motor planning processes [5, 6, 7].
-Furthermore, the work of Fathi et al.
-demonstrates that visual features located in the vicinity of a subject’s gaze during daily activities correlate with the verbal descriptions of these actions [3].
-Taken together, these insights suggest that focusing on gaze position and the associated visual features offers a pathway to understanding—and thus segmenting—object manipulations.
-In this paper, we show that a simple approach of detecting transitions between gaze landmarks during imitation enables autonomous, general, and robust task decomposition of object manipulation sequences (Fig.
+Here, the ability for an imitating agent to automatically decompose a demonstrated manipulation into multiple sub-tasks is instrumental for executing long-horizon tasks and for reusing or combining learned skills.
+In this context, a method is required that can segment demonstrations of various tasks into sub-tasks without the need for additional training or diverse data collection.
+In addition, recent imitation learning methods based on deep learning typically require hundreds of demonstrations for each task, and it is desirable to consistently decompose all of these demonstrations into the same set of sub-tasks.
+However, few methods have been proposed for such task decomposition, and none of them are able to perform it in a consistent manner.
+As a result, demonstrations are often manually segmented into sub-tasks [1], otherwise the goal state, which conditions a policy model, is usually substituted with a state a certain number of steps ahead [2, 3].
+In this study, we focus on gaze information to achieve consistent task decomposition without the need for additional training or diverse data collection.
+As Yarbus stated, “Eye movement reflects the human thought processes; so the observer’s thought may be followed to some extent from records of eye movement” [4], suggesting that gaze is a naturally evolved tool for understanding ongoing events.
+In particular, it is well established that during human object manipulations, individuals repeatedly fixate their gaze on task-relevant landmarks for certain periods and then move on to the next (Fig.
+1), and that these gaze transition behaviors are tightly coupled with the motor planning processes of hands [5, 6, 7].
+These insights suggest that leveraging an imitating agent’s gaze information, including its position and associated visual features, provides a pathway for task decomposition in demonstrated object manipulations.
+In this paper, we show that a simple approach of detecting transitions between gaze landmarks during imitation enables robust and consistent task decomposition of demonstrations (Fig.
 1).
-Here, although it would be preferable for the robot to autonomously control its own gaze during imitation, acquiring a general-purpose gaze-control policy remains challenging.
-Therefore, following the method proposed in [8], we measure the expert’s gaze during imitation (Fig.
-1) and use it as a stand-in for the robot’s gaze.
-In other words, we assume an ideal scenario of joint attention, where the imitating agent and the human demonstrator share identical gaze transitions throughout the imitation process.
-This substitution is made feasible by the fact that the expert demonstrations are generated through robot teleoperation by a human.
-Under teleoperation, object manipulation typically relies solely on visual feedback, placing the human teleoperator under the same conditions as the imitating agent, which must interpret visual information to understand the demonstrated object manipulations.
-In summary, the main contributions of this paper are follows:
-  • We propose a practical task decomposition method that decomposes object manipulations in demonstrations into sub-tasks in a simple yet robust manner, by detecting the teleoperator’s gaze transitions during demonstrations.
-  • We show that an imitation agent’s gaze control can simultaneously achieves task decomposition, underscoring the importance of developing autonomous gaze control capabilities in robotic object manipulation.
+In the current imitation learning framework, it is common to collect demonstrations via human teleoperation.
+Several approaches have been proposed that simultaneously measure the teleoperator’s gaze alongside motion recording and use it as a substitute for an imitating agent’s gaze.
+[8, 9].
+These methods improved both the robustness and dexterity of the imitated actions with minimal modifications to the conventional teleoperation system.
+Similar to these methods, we record the teleoperator’s gaze information (Fig.
+1) and use it for task decomposition.
+Our proposed method segments demonstrations at the operator’s gaze transition timings, identified using a threshold-based approach that detects abrupt changes in gaze position and visual features around the gaze point.
+To ensure consistent task decomposition across all demonstrations of the same task, we also develop a refinement process that automatically adjusts these thresholds for each demonstration.
+We apply our method to demonstration datasets of three tasks collected for imitation learning and evaluated the characteristics and consistency of the resulting sub-tasks.
+Furthermore, we investigate how the proposed refinement process improves the consistency of task decomposition and the robustness to variations in hyperparameters.
 II RELATED WORK
 II-A Robot Manipulation Imitation Using Task Decomposition
-Many approaches have been proposed that utilize temporal task decomposition for imitation learning in object manipulation tasks.
-A common strategy for enabling robustness to long-horizon tasks is to employ goal-conditioned policy models [9, 10].
+Many approaches have been proposed that utilize temporal task decomposition for imitation learning in robot manipulation.
+A common strategy for enhancing robustness in long-horizon tasks is to employ goal-conditioned policy models [1, 2].
+In these methods, the task is decomposed into multiple sub-tasks, and the goal states of each sub-task are conditioned to the model stepwise.
 Furthermore, recent years have seen an increase in language-conditioned models aimed at multi-task capabilities [10, 11, 12, 13].
-These goal- or language-conditioned methods implicitly assume that object manipulations in the given demonstrations are segmented into individual sub-tasks, and it has also been reported that providing language instructions per sub-task during inference improves task success rates [15].
+In datasets for language-conditioned models, sub-task-level instructions, such as ”pick an apple” and ”place on a plate”, are generally preferred over task-level instructions, such as ”move an apple to a plate” [14].
+These goal- or language-conditioned methods implicitly assume that object manipulations in the given demonstrations are segmented into individual sub-tasks.
 However, in most existing work, the segmentation of these demonstrations is performed manually.
-Automating this process, as proposed in our research, can enhance robot autonomy and reduce annotation costs.
+In this context, our proposed method can reduce annotation costs by automating part of the data collection process, specifically by segmenting demonstrations.
 II-B Conventional Task Decomposition Methods
-A variety of methods have been proposed for temporally segmenting actions—including object manipulation—over time.
-Nonetheless, few are suited specifically for segmenting demonstrations of robot object manipulation.
-Among the existing methods are those that segment actions based on hand or body trajectories and velocities [16, 17, 18].
-Such methods require body tracking and are not robust to trajectory variations.
-In contrast, our proposed approach does not rely on directly measuring or leveraging the motion itself.
-Other methods detect segmentation points through changes in the relationship between objects and the hand [19, 20], but they demand advanced image processing techniques, such as object or hand segmentation, detection, and tracking.
-Although further investigation is required to determine which type of segmentation is more beneficial for imitation learning, there is a qualitative difference between the segmentation results of these methods and those produced by our approach.
-Notably, our method does not necessarily cut actions at the moment of contact.
-With the advent of deep learning, there has been a rise in end-to-end approaches that take images as input and directly output action labels [21, 22, 23, 24].
-However, applying these methods to previously unseen tasks requires additional data and annotations, and they do not guarantee uniform segmentation.
-By contrast, our method can perform task decomposition without any supervision, as long as gaze data is available.
-Although there have been a few studies focusing on gaze—particularly on image features around the gaze position [25, 3]—our approach explicitly ties the timing of gaze transitions to the timing of sub-task changes, achieving a simple and uniform decomposition.
+In the field of Action Segmentation, numerous methods have been proposed for the temporal segmentation of actions and motions involving object manipulation [15, 16, 17, 18, 19, 20, 21], including ones using human gaze information [22, 23].
+However, most of these methods are primarily designed for human action recognition, and very few are suitable for task decomposition of robotic object manipulation.
+Moreover, because many of these approaches rely heavily on visual, data-driven techniques, applying them to a wide variety of robotic hardware requires the construction of new datasets tailored to each specific platform.
+Although research on task decomposition in robotic object manipulation is limited, several studies have been proposed.
+Liang et al.
+optimize the transitions between the manipulating (tool) and the manipulated (target) objects, using these transition points to segment tasks; however, their method requires the 6D poses of all relevant objects [24].
+In contrast, our proposed method does not require explicit modeling of the environment.
+Lee et al.
+reduce the dimensionality of whole-body trajectories using PCA and further model them with GMM, assigning each Gaussian component to a sub-task [25].
+In contrast, our method does not utilize hand motion data, making it robust to variations in such movements.
+Zhu et al.
+employ neural networks (NNs) for latent-space clustering to achieve task decomposition; however, their approach requires training the NN for new tasks, and the number of sub-tasks is determined solely by the time step length of each segment [26].
+In contrast, our method does not require training of NNs, thus enabling high-speed processing, and it determines the number of sub-tasks based on the number of gaze transitions, thus ensuring higher generality.
+Zhang et al.
+utilize latent vectors derived from pre-trained models specialized for object manipulation, such as R3M [27], and perform task decomposition based on the distances between these latent vectors [28].
+Similar to our method, their approach employs detection-based task decomposition without requiring additional data or training.
+However, compared to distance metrics in latent vector space, gaze transitions have been extensively studied for their coordination with hand movement planning in object manipulation, indicating that gaze-based task decomposition offers greater reliability.
 III METHOD
-In this section, we first formulate task decomposition of object manipulation based on gaze transitions.
+In this section, we first formulate the task decomposition of object manipulation based on gaze transitions.
 We then detail our implementation procedure.
-Specifically, we begin by explaining how gaze transitions are detected from time-series gaze data, and subsequently describe the refinement process used to achieve consistent segmentation across the demonstration dataset.
+Specifically, we begin by explaining how gaze transitions are detected from time-series gaze data, and subsequently describe the refinement process used to achieve consistent decomposition across the demonstration dataset.
 III-A Gaze-based Task Decomposition
-We consider a scenario in which an agent that has acquired gaze control imitates object manipulation using only visual information.
-Let $\{g_{t}\}_{t=0}^{T}$ denote the time-series of the agent’s gaze positions, and $\{f_{t}\}_{t=0}^{T}$ the time-series of the visual features observed around those gaze positions.
-When the object manipulation is decomposed into a set of sub-tasks $\{\tau_{k}\}_{k=0}^{K}$, the objective in this section is to find the $K-1$ time steps at which the sub-tasks switch (i.e.
-$\{t\mid k(t+1)=k(t)+1\}$).
-In human object manipulation, it is well-established that each task has a specific set of gaze landmarks $\{c_{i}\}_{i=0}^{N}$, and that humans fix their gaze on each landmark for a certain duration, then sequentially transition to the next landmark [6, 7, 5].
+We consider a scenario in which a human remotely controls a robot using only visual information to collect demonstrations of object manipulation for imitation learning.
+Multiple demonstrations (typically ranging from several tens to hundreds) are collected for each task, and all demonstrations of the same task share an identical set of sub-tasks.
+Let $\{g_{t}\}_{t=0}^{T}$ denote the time-series of the human teleoperator’s gaze positions, and $\{f_{t}\}_{t=0}^{T}$ the time-series of the visual features observed around those gaze positions.
+When the demonstrated task is decomposed into a set of sub-tasks $\{\tau_{k}\}_{k=0}^{K}$, the objective in this section is to determine the $K-1$ time steps at which the sub-tasks switch (i.e.
+$\{t\mid k(t+1)=k(t)+1\}$), using the gaze information $\{g_{t}\}$ and $\{f_{t}\}$.
+In human object manipulation, it is well-established that each task has a specific set of gaze landmarks $\{l_{i}\}_{i=0}^{N}$, and that humans fix their gaze on each landmark for a certain duration, then sequentially transition to the next landmark [6, 7, 5].
 Here, the landmark index $i$ is arranged according to the order in which landmarks are fixated from the start of the task (duplicates may occur).
-In this study, we hypothesize that the time steps at which the index of the fixated gaze landmark changes, $\{t\mid i(t+1)=i(t)+1\}$, corresponds to the time steps at which the sub-tasks switch, $\{t\mid k(t+1)=k(t)+1\}$.
+In this study, we hypothesize that the time steps at which the index of the fixated gaze landmark changes, $\{t\mid i(t+1)=i(t)+1\}$, correspond to the time steps at which the sub-tasks switch, $\{t\mid k(t+1)=k(t)+1\}$.
 In other words, we realize task decomposition by detecting transitions of gaze landmarks.
-Let $p^{i(t)}_{t}$ be the location of the landmark $c_{i(t)}$ in the agent’s field of view at time step $t$.
+Let $p^{i(t)}_{t}$ be the location of the landmark $c_{i(t)}$ in the teleoperator’s field of view at time step $t$.
 Then, the gaze position $g_{t}$ and the visual feature $f_{t}$ at that gaze position can be written as follows:
-$g_{t}=p\bigl{(}c_{i=i(t)},t\bigr{)}+\epsilon_{t},$ (1)
-$f_{t}=f\bigl{(}g_{t}\bigr{)}+\zeta_{t}=f\bigl{(}p(c_{i=i(t)},t)+\epsilon_{t}\bigr{)}+\zeta_{t}.$ (2)
-where $\epsilon_{t}$ and $\zeta_{t}$ represent noise in the gaze position and perturbations in the visual features caused, for example, by hand movements, respectively, and $f(\cdot)$ is the visual feature extractor at the given gaze position.
+$g_{t}=p^{i(t)}_{t}+\epsilon_{t},$ (1)
+$f_{t}=f\bigl{(}g_{t}\bigr{)}+\zeta_{t}=f\bigl{(}p^{i(t)}_{t}+\epsilon_{t}\bigr{)}+\zeta_{t},$ (2)
+where $\epsilon_{t}$ and $\zeta_{t}$ represent noise in the gaze position and perturbations in the visual features caused, for example, by the end-effector movements, respectively, and $f(\cdot)$ is the visual feature extractor at the given gaze position.
 In other words, by leveraging the relationships in Eqs.
-(1) and (1), we can detect the transitions of gaze landmarks from the gaze information $\{g_{t}\}$ and $\{f_{t}\}$, and thereby achieve task decomposition of the demonstrated object manipulation.
+1 and 2, we can detect the transitions of gaze landmarks $\{l_{i}\}$ from the gaze information $\{g_{t}\}$ and $\{f_{t}\}$, and thereby achieve task decomposition of the demonstrated object manipulation.
 III-B Gaze Transition Detection
-We now describe the implementation details for segmenting demonstration tasks based on gaze transitions.
-As discussed in the previous section, an algorithm for detecting the time steps at which gaze landmarks switch, using time-series gaze data, is critical for task decomposition.
+We now describe the implementation details for segmenting demonstrations based on gaze transitions.
+As discussed in the previous section, an algorithm for detecting the time steps at which gaze landmarks switch, using time-series gaze data, is critical for the task decomposition.
 III-B1 Median Filtering
 Gaze data typically contain noise.
-Even when the agent is fixating on a stationary gaze landmark, the recorded gaze positions often exhibit fluctuations.
-Moreover, the agent may occasionally shift its gaze for a short period to a different landmark during object manipulation.
-To handle such noise in a simple yet effective manner, we adopted a median-filtering approach.
-Specifically, for the agent’s gaze position $g_{t}$ we applied the process shown in Eq.
-(3) with a window size $w$ as a hyperparameter, and used the filtered gaze data $\hat{g}_{t}$ in all subsequent processing:
+Even when the teleoperator is fixating on a stationary gaze landmark, the recorded gaze positions often exhibit fluctuations.
+Moreover, the teleoperator may occasionally shift its gaze for a short period to a different landmark during object manipulation.
+To handle such noise in a simple yet effective manner, we adopt a median-filtering approach.
+Specifically, for the teleoperator’s gaze position $g_{t}$, we apply the process shown in Eq.
+3 with a window size $w$ as a hyperparameter, and use the filtered gaze data $\hat{g}_{t}$ in all subsequent processing:
 $\hat{g}_{t}=\text{median}\bigl{(}g_{t-\tfrac{w}{2}:t+\tfrac{w}{2}}\bigr{)}.$ (3)
@@
 Gaze transitions cause not only positional changes in the gaze but also variations in the visual features around the gaze.
-To detect these transitions from visual features, it is necessary to employ a feature extractor based on high-quality representation learning.
-For each time step, we cropped an $b\times b$ region centered at the filtered gaze position $\hat{g}_{t}$ from both the left and right images and fed these cropped images into the pre-trained CLIP model [26] to extract their feature vectors $f_{t}\in\mathbb{R}^{512\times 2}$.
-This time-series of feature vectors, along with the gaze position, was used for gaze transition detection.
+We also consider detecting gaze transitions using visual features.
+To robustly detect changes in visual features, it is essential to employ a feature extractor based on high-quality representation learning.
+For each time step, we crop an $b\times b$ region centered at the filtered gaze position $\hat{g}_{t}$ from both the left and right images and feed these cropped images into the pre-trained CLIP model [29] to extract their feature vectors $f_{t}\in\mathbb{R}^{512\times 2}$.
+This time-series of feature vectors, along with the gaze position, is used for gaze transition detection in this study.
 III-B3 Change Detection
-For the obtained gaze position $\hat{g}_{t}$ and visual feature $f_{t}$, we computed the difference from the previous time step to detect abrupt changes in the time-series data caused by gaze transitions:
+For the obtained gaze position $\hat{g}_{t}$ and visual feature $f_{t}$, we compute the difference from the previous time step to detect abrupt changes in the time-series data caused by gaze transitions:
 $s^{pos}_{t}=\bigl{\lVert}\hat{g}_{t}-\hat{g}_{t-1}\bigr{\rVert},$ (4)
 $s^{feat}_{t}=-\frac{\log\bigl{(}{f^{left}_{t-1}}^{T}f^{left}_{t}+1\bigr{)}+\log\bigl{(}{f^{right}_{t-1}}^{T}f^{right}_{t}+1\bigr{)}}{2}$
-$+\log 2.$ (5)
-Here, $s^{pos}_{t}$ and $s^{feat}_{t}$ represent the change scores for the gaze position and the visual features, respectively.
+$+\log 2,$ (5)
+where $s^{pos}_{t}$ and $s^{feat}_{t}$ represent the change scores for the gaze position and the visual features, respectively.
 Specifically, $s^{pos}_{t}$ is the Euclidean distance between the two 4-dimensional vectors representing the left and right gaze positions.
 Meanwhile, $s^{feat}_{t}$ is derived from the inner products between the corresponding gaze feature vectors in consecutive time steps, normalized to ensure non-negativity, and then averaged for the left and right features.
-We introduced the thresholds $\theta_{pos}$ and $\theta_{feat}$ as hyperparameters for these two change scores, and identified any time step at which both change scores exceeded their respective thresholds as a change point $d_{i}\in D$.
-III-C Refinement Based on Segmentation Uniformity
-From a machine learning perspective—particularly one that relies on multiple demonstrations—it is desirable for all demonstrations of the same task to undergo a uniform task decomposition.
-For example, in a pick-and-place task, mixing one demonstration segmented into “pick” and “place” with another in which these two motions are grouped into a single segment is not ideal for building a consistent training dataset.
-Therefore, in this study, we introduce a refinement procedure, $D\rightarrow D^{*}$, for each demonstration to ensure that demonstrations of the same task share as consistent a number of sub-tasks as possible.
-First, we enumerate the number of change points $\lvert D\rvert$ obtained via the gaze transition detection (Section III-B) for all demonstrations and denote the most frequent value by $s$.
+We introduce the thresholds $\theta_{pos}$ and $\theta_{feat}$ as hyperparameters for these two change scores, and identify any time step at which both change scores exceed their respective thresholds as a change point $c_{i}\in C$.
+III-C Refinement Based on Decomposition Consistency
+In recent machine learning–based imitation learning methods, particularly those relying on multiple demonstrations, all demonstrations of the same task typically comprise a consistent set of sub-tasks.
+In such scenarios, it becomes possible to further improve the consistency of task decomposition across all these demonstrations.
+Here, we introduce a refinement procedure, $C\rightarrow C^{*}$, to ensure that all demonstrations of the same task share as consistent a number of sub-tasks as possible.
+The algorithm for this process is summarized in Algorithm 1.
+First, we enumerate the number of change points $\lvert C\rvert$ obtained via the gaze transition detection (Section III-B) for every demonstration and denote the most frequent value by $s$.
 We assume $s$ to be the number of sub-tasks for the task and adjust the thresholds $\theta_{pos}$ and $\theta_{feat}$ so that the number of segments in each demonstration matches $s$.
@@
 If no threshold values are found that yield $s$ segments within a maximum number of iterations, that demonstration is treated as a detection failure and excluded from the training data.
-Finally, the refined set of change points $D^{*}$ obtained by this process is taken as the sub-task boundary for each demonstration, enabling us to divide as many demonstrations as possible into a consistent number of sub-tasks.
+Finally, the refined set of change points $C^{*}$ obtained by this process is taken as the sub-task boundary for each demonstration, enabling us to divide as many demonstrations as possible into a consistent number of sub-tasks.
+Algorithm 1 Task Decomposition Refinement Process
+1:Given: A set of demonstrations for the same task:
+$D=\{(s_{t}^{\mathrm{pos}},\,s_{t}^{\mathrm{feat}})^{(n)}\}_{n=1}^{N},\quad t\in[0,\,T^{(n)}],$
+where $N$ is the number of demonstrations and $T^{(n)}$ is the total time steps in the $n$-th demonstration.
+2:Notation: Let the function $\textit{detect}(\cdot)$ perform change point detection:
+$C\;\leftarrow\;\textit{detect}\bigl{(}\{s_{t}^{\mathrm{pos}}\},\,\{s_{t}^{\mathrm{feat}}\},\,\theta_{\mathrm{pos}},\,\theta_{\mathrm{feat}}\bigr{)}.$
+3:Initialize: $\theta_{\mathrm{pos}},\theta_{\mathrm{feat}}\;\leftarrow\;\text{default values}$
+4:for $n=1$ to $N$ do
+5:     $C^{(n)}\;\leftarrow\;\textit{detect}\bigl{(}\{s_{t}^{\mathrm{pos}}\}^{(n)},\,\{s_{t}^{\mathrm{feat}}\}^{(n)},\,\theta_{\mathrm{pos}},\,\theta_{\mathrm{feat}}\bigr{)}$
+6:end for
+7:$s\;\leftarrow\;\text{mode}(\{|C^{(n)}|\}_{n=1}^{N})$
+8:for $n=1$ to $N$ do
+9:     while $\bigl{|}C^{(n)}\bigr{|}<s$ do
+[…]
+11:         $C^{(n)}\leftarrow\textit{detect}(\{s_{t}^{\mathrm{pos}}\}^{(n)},\{s_{t}^{\mathrm{feat}}\}^{(n)},\theta_{\mathrm{pos}},\theta_{\mathrm{feat}})$
+12:     end while
+13:     while $\bigl{|}C^{(n)}\bigr{|}>s$ do
+[…]
+15:         $C^{(n)}\leftarrow\textit{detect}(\{s_{t}^{\mathrm{pos}}\}^{(n)},\{s_{t}^{\mathrm{feat}}\}^{(n)},\theta_{\mathrm{pos}},\theta_{\mathrm{feat}})$
+16:     end while
+17:     $C^{*(n)}\;\leftarrow\;C^{(n)}$
+18:end for
+19:Return: $\{C^{*(n)}\}_{n=1}^{N}$
 IV EXPERIMENTS
 IV-A Robot System
-Our group previously developed a dual-arm robotic system designed for imitation learning through teleoperation [8].
-A human teleoperator controls the robot remotely by viewing its environment through a head-mounted display (HMD), which simultaneously records the operator’s eye gaze.
-The system can be configured to use either a physical dual-arm robot (a pair of UR5 arms) or a simulated counterpart, enabling the collection of expert demonstrations in both real and virtual settings [27].
-The recorded gaze data are provided as pixel coordinates corresponding to the images displayed on the HMD.
-In this study, we evaluated our task decomposition method using an expert demonstration dataset collected by the authors.
-Each demonstration includes synchronized left and right image streams, along with corresponding gaze data, and all of these were employed in our experiments.
+We used a dual-arm robotic system that was previously developed for imitation learning through teleoperation [8].
+A human teleoperator controlled the robot remotely by viewing its environment through a head-mounted display (HMD), which simultaneously recorded the operator’s eye gaze.
+The system can be configured to use either a physical dual-arm robot, a pair of UR5 robots from Universal Robots Inc., or a simulated counterpart, enabling the collection of expert demonstrations in both real and virtual settings [30].
+The recorded gaze data were provided as pixel coordinates corresponding to the images displayed on the HMD.
+In this study, we evaluated our task decomposition method using an expert demonstration dataset collected through this system.
+Each demonstration included synchronized left and right image streams, along with corresponding gaze data, and all of these were employed in our experiments.
 IV-B Task Setup
@@
 We first collected approximately 100 demonstrations for each task as follows:
-  • WrapCandy (real robot): In this task, a green candy and a brown handkerchief are arranged on a table with a green tablecloth.
+  • WrapCandy (real robot, 64 demos): In this task, a green candy and a brown handkerchief are arranged on a table with a green tablecloth.
 The position and orientation of the two objects are randomized on the table.
@@
 The robot then picks up the (top) left corner of the handkerchief and folds it to wrap the candy.
-  • PenInCup (real robot): In this task, a blue marker pen and a red cup are placed on a table; the position and orientation of the two objects is randomized on the table.
+  • PenInCup (real robot, 125 demos): In this task, a blue marker pen and a red cup are placed on a table; the position and orientation of the two objects is randomized on the table.
 The robot picks up the pen and puts it in the cup.
-  • PileBox (simulator robot): In this task, red and blue boxes are placed on a wooden table; the position and orientation of the two boxes is randomized.
+  • PileBox (simulated robot, 147 demos): In this task, red and blue boxes are placed on a wooden table; the position and orientation of the two boxes is randomized.
 The robot grabs the red box and places it on top of the green box.
@@
 Fig.
-2 presents an representative example of segmentation results for each task.
+2 presents a representative example of segmentation results for each task.
 Figure 2: Typical task decomposition results using the proposed method for the tasks of WrapCandy (a), PenInCup (b), and PileBox (c).
@@
 A closer examination of the states before and after each segmentation revealed two common features across the tasks:
-  • Each sub-task began when the agent (released the previous object and) started reaching for the next relevant object.
+  • Each sub-task began when the robot (released the previous object and) started reaching for the next relevant object.
   • Each sub-task ended once the action’s completion was visually confirmed.
 (In a grasping action, for example, the sub-task did not end when the hand merely contacted the object, but rather when the object was lifted slightly off the table, confirming a successful grasp.)
-Crucial aspects of object manipulation exist even before actual contact with the object.
-For example, how the agent reaches for the object significantly influences the success of the manipulation.
-For this reason, non-contact actions such as reaching should also be included in the segment.
-However, methods that rely solely on outcome-based cues, such as contact events, cannot detect these reaching actions.
-IV-D Uniformity of Task Decomposition
-From a practical perspective, it is desirable that all demonstrations within each task exhibit a uniform pattern of task decomposition.
+IV-D Consistency of Task Decomposition
+In this section, we evaluated the consistency of the resulting task decomposition.
 Table I shows the number of demonstrations classified as “minority data,” in which the segmentation timings or the number of sub-tasks differed from the most common (majority) decomposition for each task.
-In addition, we also evaluated decomposition uniformity under the following ablation settings: (1) without the refinement process (w/o refine), (2) detecting changes based solely on the gaze position $\hat{g}_{t}$ (w/o feat), (3) detecting changes based solely on visual features around the gaze $f_{t}$ (w/o pos), and (4) omitting the refinement process from these two settings (w/o feat, refine and w/o pos, refine).
+In addition to the proposed approach, we also evaluated the decomposition consistency under the following ablation settings: (1) without the refinement process (w/o refine), (2) detecting changes based solely on the gaze position $\hat{g}_{t}$ (w/o feat), (3) detecting changes based solely on visual features around the gaze $f_{t}$ (w/o pos), and (4) omitting the refinement process from these two settings (w/o feat, refine and w/o pos, refine).
 TABLE I: Number of minority demonstrations exhibiting task decomposition with different timing or segment counts from the majority in each task.
@@
 - w/o pos, refine 4 4 45 53
-When the refinement process was applied, we found that all demonstrations across all tasks were segmented in a consistent manner.
+As illustrated in Table I, we found that all demonstrations across all tasks were segmented in a consistent manner when the refinement process was applied.
 In contrast, removing the refinement process led to some demonstrations whose decomposition differed from the majority.
@@
 In each scenario, we observed instances of over-segmentation when the gaze shifted slightly during an action, or missed segmentation when the gaze displacement was too small (e.g., when objects were close together).
-Furthermore, in the condition where only gaze features were employed, change detection was triggered whenever the hand or an object grasped by the hand entered the cropped image region, resulting in over-segmentation.
+Furthermore, in the condition where only gaze features were employed, change detection was triggered whenever the end-effector or a grasped object entered the cropped image region, resulting in over-segmentation.
 Figure 3: Number of successfully decomposed demonstrations when varying the hyperparameters. (a) Without refinement process. (b) With refinement process.
@@
 Specifically, $w$ determines the filtering strength of the median filter applied to the gaze position, while $\theta_{pos}$ determines the sensitivity for detecting gaze transitions.
-Figure 3 shows the number of demonstrations that were correctly segmented when $w$ was set to $\{5,10,15,20,25,30\}$ and $\theta_{pos}$ was set to $\{25,50,75,100,125,150\}$.
-When hyperparameters varied extensively, the number of demonstrations that resulted in incorrect decomposition exhibited gradual fluctuations in the absence of refinement processing.
-In contrast, with refinement processing applied, decomposition was successfully achieved for all demonstrations across the broad parameter range.
+Fig.
+3 shows the number of demonstrations that were correctly segmented when $w$ was set to $\{5,10,15,20,25,30\}$ and $\theta_{pos}$ was set to $\{25,50,75,100,125,150\}$.
+When hyperparameters varied extensively, the number of demonstrations that resulted in incorrect decomposition exhibited gradual fluctuations in the absence of refinement processing (Fig.
+3).
+In contrast, with refinement processing applied, decomposition was successfully achieved for all demonstrations across the broad parameter range (Fig.
+3).
 These findings allow us to conclude that the proposed method possesses the robustness necessary for application to various robotic systems.
-In most cases, incorrect decomposition persisted after the refinement process (including instances where the number of successfully segmented demonstrations decreased post-refinement) because decomposition errors resulting from an incorrect number of segments had already become predominant beforehand.
+In cases where incorrect task decomposition persisted after the refinement process, including instances where the number of successfully segmented demonstrations was lower with refinement (Fig.
+3) than without it (Fig.
+3), decomposition errors resulting from an incorrect number of segments had already become predominant beforehand.
 Consequently, instead of achieving its intended purpose, the refinement process altered demonstrations that had been correctly segmented initially, leading to an incorrect number of sub-tasks.
 Furthermore, setting the window width $w$ to 25 or higher caused the influence of median filtering to become excessively strong.
-As a result, particularly in the WrapCandy task, it became impossible to detect the required number of segmentation points during the refinement process, regardless of how much the threshold $\theta_{pos}$ was lowered.
+As a result, it became impossible to detect the required number of segmentation points during the refinement process, regardless of how much the threshold $\theta_{pos}$ was lowered.
 V CONCLUSIONS
-In this study, we proposed a robust and generalizable method for the autonomous task decomposition of object manipulation behaviors, which serves as a foundation for enabling robots, as autonomous agents, to move beyond rote imitation and to understand appropriately demonstrated object manipulations.
-This method achieves segmentation solely through detecting the agent’s gaze transitions.
-Notably, the incorporation of a refinement process allows for consistent task decomposition, which is particularly desirable in applications such as machine learning.
-Our findings suggest that gaze plays a crucial role in cognition and learning during object manipulation for autonomous agents, including humans.
-Once a robot acquires gaze control, it can autonomously perform human-like segmentation, which serves as the basis for diverse learning processes, relying solely on the observation of its own gaze transitions.
+In this study, we proposed a practical method that provides a consistent and robust task decomposition for robot manipulation in a recent imitation learning framework.
+The most important finding is that the desired task decomposition was successfully achieved for all demonstrations by adjusting the threshold in the refinement process.
+This result suggests that utilizing gaze transitions for task decomposition is fundamentally effective.
+However, the proposed method has several limitations.
+Based on the characteristics of the current imitation learning dataset, we assumed that all demonstrations for a given task consist of the same set of sub-tasks.
+In contrast, in natural settings, the actions taken to accomplish the same task may consist of a different number of sub-tasks.
+For example, when pouring water into a cup, one may either grasp the cup before pouring or pour without grasping it.
+In even more general settings, demonstrations are no longer labeled by task, and a single demonstration may include multiple tasks, as seen in play data [31].
+In such cases, our gaze transition detection can be performed, but the refinement process cannot be applied, resulting in several percent incorrect task decomposition.
 To achieve greater autonomy in the future, robots must be capable of directly observing and imitating human behaviors rather than relying on teleoperation.
-In such situations, even if human gaze data were available, using it as a substitute for the robot’s gaze would be inappropriate because humans utilize other sensory modalities, such as tactile information, in addition to visual input for object manipulation.
-Therefore, breakthroughs in autonomous gaze control are urgently needed today, either for such cases or for cases where demonstration datasets simply lack gaze data from teleoperators, in order to make robots more intelligent.
+In such a scenario, the robot can no longer use the expert’s gaze as supervision.
+Therefore, breakthroughs in autonomous gaze control are urgently needed so that the robot can determine where to focus during the demonstration in an unsupervised manner.